# The 3D-PC: a benchmark for visual perspective taking in humans and machines

### v1

@@ 5 Discussion
A key limitation of our study is that our version of VPT represents the most basic form studied in the developmental psychology literature. While solving this task is evidently an extraordinary challenge for DNNs, it is only one small step towards human-level capabilities for reasoning about 3D worlds in general. Far more research is needed to identify additional challenges, architectures, and training routines that can help DNNs perceive and reason about the world like humans do. We release our 3D-PC data and code at https://github.com/serre-lab/VPT to support this goal.
6 Acknowledgements
@@
Funding for this project was provided by the Office of Naval Research (N00014-19- 1-2029) and ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A0004). Additional support was provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] In the discussion.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] Appendix section A.3.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See methods.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Appendix section A.5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We report error bars over human performance in all figures. We also report model/error bars in performance and correlation with humans (Appendix fig. A.3).
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Appendix section A.5.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We used existing models and libraries and cited each.
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix section A.1.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See methods.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See methods.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See appendix section A.8
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See methods.
@@ A.2 Data Access
We release benchmarking code and data download instructions at https://github.com/serre-lab/VPT.
@@ A.4 Data Generation
To generate data for the 3D-PC, we first trained 3D Gaussian Splatting [7] models on videos from the Common Objects in 3D (Co3D) [43], which yielded 3D representations of each scene. We then imported trained models into Unity [44] using Unity Gaussian Splatting [45] and added 3D models of the green camera and red ball to each. Finally, we rendered 50 images along a smooth viewpoint camera trajectory sampled near the original trajectory used for training the Gaussian Splatting model. For each 3D scene, we created 5 positive and 5 negative settings for VPT.
@@ A.5 Model Zoo
We linearly probed 317 DNNs from Pytorch Image Models (TIMM) [46] (Table 1) along with foundational vision models following the procedures in  [23]. All DNNs were trained and evaluated with NVIDIA-RTX 3090 GPUs from the Brown University Center for Computation & Visualization. All linear probes were trained for 50 epochs, with a $45e-$ 4 learning rate, a $41e-$ 4 weight decay, a $0.3$ dropout rate, and a batch size of 128. We fine-tuned each of the TIMM models for 30 epochs, a $55e-$ 5 learning rate, $41e-$ 4 weight decay, $0.7$ dropout rate, and a batch size of 16. Linear probing took approximately 20 minutes per model, and fine-tuning varied from 3 to 24 hours on a NVIDIA-RTX 3090 GPU.
@@
CNN ConvMixer [61] 3
ConvNeXT [62] 10
DenseNet [63] 4
DLA [64] 5
DPN [65] 6
EfficientNet [66] 4
GhostNet [67] 1
HRNet [68] 8
LCNet [69] 3
MixNet [70] 4
MnasNet [71] 3
MobileNet [72] 14
RegNet [73] 6
Res2Net [74] 5
ResNet [75] 26
ResNeSt [76] 3
RexNet [77] 5
ResNext [78] 2
SPNASNet [79] 1
TinyNet [80] 2
VGG [81] 14
Transformer BEiT [82] 9
CAFormer [83] 6
CaiT [84] 3
ConViT [85] 3
CrossViT [86] 2
DaViT [87] 3
DeiT [88] 12
EfficientFormer [89] 7
EVA [48] 9
FocalNet [90] 6
LeViT [91] 5
MaxViT [92] 6
MobileViT [93] 3
MViT [94] 3
PiT [95] 8
PVT [96] 7
Swin [60] 16
Twins-SVT [97] 5
ViT [98] 36
Volo [99] 7
XCiT [100] 6
PoolFormer [101] 8
Hybrid CoaT [102] 7
CoAtNet [103] 8
EdgeNeXt [104] 1
Visformer [105] 2
Foundation Depth Anything [53] 1
DINOv2 [50] 1
iBoT [51] 1
MAE [49] 1
MiDas [15] 1
SAM [52] 1
VLM ChatGPT4 [27] 1
Gemini [54] 1
Claude 3 [55] 1
Diffusion Stable Diffusion 2.0 [28] 1
@@ A.6 VLM Evaluation
The prompt we used for the depth task was “In this image, is the red ball closer to the observer or is the green arrow closer to the observer? Answer only BALL if the red ball is closer, or ARROW if the green arrow is closer, nothing else.” and the prompt for the VPT-basic task was “In this image, if viewed from the perspective of the green 3D arrow in the direction the arrow is pointing, can a human see the red ball? Answer only YES or NO, nothing else”. We evaluated each model’s generated responses across multiple temperatures, ranging from $0.0$ to $0.7$ in increments of $0.1$, and we report the average of the best 3 runs. Note that while this evaluation approach gives the VLMs more opportunities to perform well on our benchmark than other models, they still struggled immensely (see main text).
A.7 Stable Diffusion Evaluation
We followed the method of Li et al. [14] to evaluate Stable Diffusion 2.0 on the 3D-PC. This involved trying multiple prompts to optimize the zero-shot classification performance of the Stable Diffusion 2.0 model, on VPT-basic and depth order tasks. For VPT-basic we found that the prompt "A photo with red ball is visible from the green arrow’s perspective" for positive class and "A photo with red ball not visible from the green arrow’s perspective" for the negative class led to the best performance. For the depth order task, the prompt with the highest performance was "A photo with green arrow closer to the camera as compared to red ball" and "A photo with red ball closer to the camera as compared to green arrow" for positive and negative classes respectively.
A.8 Human Benchmark
@@
A.8.1 Experiment design
At the outset of the experiment, we acquired participant consent through a form approved by the Brown University’s Institutional Review Board (IRB). The experiment was performed on a computer using the Chrome browser. Following consent, we presented a demonstration with instructions and an example video. Participants had the option to revisit the instructions at any time during the experiment by clicking a link in the top right corner of the navigation bar.
Figure A.1: An experiment trial.
Figure A.2: The consent screen.
@@ A.8.1 Experiment design
A.9 Human vs. DNN decision making on VPT-basic
@@
We compared the decision strategies of humans and DNNs on VPT-basic by measuring the correlations between their error patterns with Cohen’s $\kappa$ [106]. Model $\kappa$ scores were mostly correlated with accuracy on VPT-basic after linear probes and fine-tuning (Fig. A.3). However, while nearly all DNNs were highly correlated with human error patterns after fine-tuning, the correlation between $\kappa$ scores and task accuracy disappeared (Fig. A.3B, purple dots).
@@
Figure A.3: Error pattern correlations (Cohen’s $\kappa$) between humans and DNNs on VPT-basic
A.10 Datasheet for datasets
Motivation
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
The dataset was designed to test 3D perception in humans and DNNs, with an emphasis on the capabilities of each for visual perspective taking (VPT). Humans rely on VPT everyday for navigating and socializing, but despite its importance, there has yet to be a systematic evaluation of this ability in DNNs.
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
This dataset was created by this paper authors, who are affiliated with the Carney Institute for Brain Science at Brown University and the Cognitive Sciences Department at UC Irvine.
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
Funding for this project was provided by the Office of Naval Research (N00014-19- 1-2029) and ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A0004). Additional support provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
The instances contain images of real-world objects and scenes along with shapes generated with computer graphics.
How many instances are there in total (of each type, if appropriate)?
There are 7574 images in the training and testing sets.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
We release all data.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
Each instance consists of an image rendered from 3d Gaussian Splatting [7] models trained on Co3D [43] scenes.
Is there a label or target associated with each instance? If so, please provide a description.
The images are labeled for VPT and depth order tasks. In the VPT task, an image is labeled as positive when the red ball is visible from the green camera’s perspective. In the depth task, an image is labeled as positive when the red ball is further away than the green arrow from the viewer. For both tasks, we label positives as 1 and negatives as 0.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
N/A
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
N/A
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
We provide training, validation and testing splits in the released dataset. The training set contains images rendered from 20 unique scenes from 10 categories. The testing set images are rendered from 10 additional scenes from the same categories. We randomly selected 10% of the training set as the validation set.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
N/A
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
The dataset uses videos from the Co3D dataset [43], which is publicly available under CC BY-NC 4.0 license.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals non-public communications)? If so, please provide a description.
N/A
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
N/A
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
Yes
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
No
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
No, all results are anonymous.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
N/A
Any other comments?
Collection Process
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
All images were rendered from 3D gaussian splatting [7] models trained on videos from Co3D [43]. We imported the model into Unity [44, 45] to render images.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
We used Unity [44] and Unity Gaussian Splatting [45] to edit the scenes and label them in 3D view.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
N/A
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
The paper’s authors were involved in the data collection process.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
N/A
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
Yes
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
As described in the Methods, we collected data from online participants through Prolific, and we also collected data in-person for several subjects.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
Yes. See Section A.8 for details.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Yes. See Fig A.2 for the consent screen with the exact language used.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
Yes. The participants were provided with our contact information and were encouraged to reach out in such cases.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
Our experiment was approved by the IRB board at Brown University.
Any other comments?
Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
We used Unity to label images for VPT and depth tasks. We removed images where the objects of interest (red ball and green camera) were not visible.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
N/A
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
N/A
Any other comments?
Uses
Has the dataset been used for any tasks already? If so, please provide a description.
We evaluated vision DNNs on the dataset. Please refer to the main paper for details.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
The code and data are publicly available at https://github.com/serre-lab/VPT
What (other) tasks could the dataset be used for?
We mainly expect the dataset to be used for evaluating 3D perception capabilities of new vision or vision-language DNNs.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
N/A
Are there tasks for which the dataset should not be used? If so, please provide a description.
N/A
Any other comments?
Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
Yes, we will release the dataset to the public at https://github.com/serre-lab/VPT
How will the dataset be distributed (e.g., tarball on website, API, GitHub) Does the dataset have a digital object identifier (DOI)?
We provide download instructions at https://github.com/serre-lab/VPT
When will the dataset be distributed?
The dataset is available from June 5th, 2024.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
We release our data under a Creative Commons CC-BY license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
N/A
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
N/A
Any other comments?
Maintenance
Who will be supporting/hosting/maintaining the dataset?
The authors will be hosting and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Contact the corresponding author through email.
Is there an erratum? If so, please provide a link or other access point.
N/A
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
We are actively working on expanding the dataset with new instances and tasks. We will update our GitHub repository accordingly for any dataset update.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
Human participant data was de-identified, and there are no time limits on its retention.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Yes, we will maintain old versions of the dataset on our website.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
We are open to any suggestions and contributions through our GitHub repository. https://github.com/serre-lab/VPT

### v2

@@ 5 Discussion
Limitations
One key limitation of our study is that we were not able to isolate the precise cues that humans versus DNNs use to solve depth order. While both were very capable of solving the task, it’s very possible that both relied on different monocular depth cues. For example, the relative size of objects is a very salient cue for solving the task. However, object occlusion and interposition is another strong cue in this dataset, and it’s unclear if DNNs and humans rely on one or the other, or both, or other cues like the lighting and shadows of objects. Future work that can drill down on the strategies of humans and DNNs for perceiving 3D spatial properties will make significant contributions for understanding the alignment of biological and machine vision.
@@
Another key limitation of our study is that our version of VPT represents the most basic form studied in the developmental psychology literature. While solving this task is evidently an extraordinary challenge for DNNs, it is only one small step towards human-level capabilities for reasoning about 3D worlds in general. Far more research is needed to identify additional challenges, architectures, and training routines that can help DNNs perceive and reason about the world like humans do. We release our 3D-PC data, 3D Gaussian Splatting models, and code anonymously at https://huggingface.co/datasets/3D-PC/3D-PC to support this goal.
Finally, we validated developmental psychology work that humans rely on a line-of-sight strategy to solve VPT, and demonstrated in our VPT-Strategy task that DNNs learn to use a different strategy after fine-tuning. We provided evidence through our attribution map experiment (Fig. 10) that DNN strategies are highly focused on the target objects. However, we were not able to characterize exactly how DNNs use object features for VPT, which may be critical for developing the next-generation of vision models that can solve VPT like humans do.
Acknowledgments
@@
Our work is supported by ONR (N00014-24-1-2026, NSF (IIS-2402875), the ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A-0004). Additional support was provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
@@ A.2 Data Access
We release benchmarking code, data, and 3D Gaussian Splatting models anonymously at https://huggingface.co/datasets/3D-PC/3D-PC.
@@ A.4 Data Generation
To generate data for the 3D-PC, we first trained 3D Gaussian Splatting (Kerbl et al., 2023) models on videos from the Common Objects in 3D (Co3D) (Reizenstein et al., 2021), which yielded 3D representations of each scene. We then imported trained models into Unity (Juliani et al., 2018) using Unity Gaussian Splatting (Pranckevicius, 2023) and added 3D models of the green camera and red ball to each. Finally, we rendered 50 images along a smooth viewpoint camera trajectory sampled near the original trajectory used for training the Gaussian Splatting model. For each 3D scene, we created 5 positive and 5 negative settings for VPT.
@@ A.5 Model Zoo
We linearly probed 317 DNNs from Pytorch Image Models (TIMM) (Wightman, 2019) (Table 1) along with foundational vision models following the procedures in  (El Banani et al., 2024). All DNNs were trained and evaluated with NVIDIA-RTX 3090 GPUs. All linear probes were trained for 50 epochs, with a $45e-$ 4 learning rate, a $41e-$ 4 weight decay, a $0.3$ dropout rate, and a batch size of 128. We fine-tuned each of the TIMM models for 30 epochs, a $55e-$ 5 learning rate, $41e-$ 4 weight decay, $0.7$ dropout rate, and a batch size of 16. Linear probing took approximately 20 minutes per model, and fine-tuning varied from 3 to 24 hours on a NVIDIA-RTX 3090 GPU.
@@
CNN AlexNet (Krizhevsky et al., 2012b) 1
ConvMixer (Trockman & Kolter, 2022) 3
ConvNeXT (Liu et al., 2022) 10
DenseNet (Huang et al., 2018) 4
DLA (Yu et al., 2019) 5
DPN (Chen et al., 2017) 6
EfficientNet (Tan & Le, 2020) 4
GhostNet (Han et al., 2020b) 1
HRNet (Sun et al., 2019) 8
LCNet (Cui et al., 2021) 3
MixNet (Tan & Le, 2019) 4
MnasNet (Tan et al., 2019) 3
MobileNet (Howard et al., 2019) 14
RegNet (Radosavovic et al., 2020) 6
Res2Net (Gao et al., 2021) 5
ResNet (He et al., 2015) 26
ResNeSt (Zhang et al., 2020) 3
RexNet (Han et al., 2020a) 5
ResNext (Xie et al., 2016) 2
SPNASNet (Stamoulis et al., 2019) 1
TinyNet (Han et al., 2020c) 2
VGG (Simonyan & Zisserman, 2014) 14
Transformer BEiT (Bao et al., 2021b) 9
CAFormer (Yu et al., 2022b) 6
CaiT (Touvron et al., 2021b) 3
ConViT (d’Ascoli et al., 2021) 3
CrossViT (Chen et al., 2021a) 2
DaViT (Ding et al., 2022) 3
DeiT (Touvron et al., 2021a) 12
EfficientFormer (Li et al., 2022b) 7
EVA (Fang et al., 2023) 9
FocalNet (Yang et al., 2022) 6
LeViT (Graham et al., 2021) 5
MaxViT (Tu et al., 2022) 6
MobileViT (Mehta & Rastegari, 2022) 3
MViT (Li et al., 2022a) 3
PiT (Heo et al., 2021) 8
PVT (Wang et al., 2022) 7
Swin (Liu et al., 2021) 16
Twins-SVT (Chu et al., 2021) 5
ViT (Dosovitskiy et al., 2021b) 36
Volo (Yuan et al., 2022) 7
XCiT (El-Nouby et al., 2021) 6
PoolFormer (Yu et al., 2022a) 8
Hybrid CoaT (Xu et al., 2021) 7
CoAtNet (Dai et al., 2021) 8
EdgeNeXt (Maaz et al., 2022) 1
Visformer (Chen et al., 2021b) 2
Foundation Depth Anything (Yang et al., 2024) 1
DINOv2 (Oquab et al., 2023) 1
iBoT (Zhou et al., 2021) 1
MAE (He et al., 2022) 1
MiDas (Ranftl et al., 2022) 1
SAM (Kirillov et al., 2023) 1
VLM ChatGPT4 (Achiam et al., 2023) 1
Gemini (Team et al., 2023) 1
Claude 3 (Anthropic, 2024) 1
Diffusion Stable Diffusion 2.0 (Rombach et al., 2021) 1
@@ A.6 VLM Evaluation
The prompt we used for the depth task was ‘‘In this image, is the red ball closer to the observer or is the green arrow closer to the observer? Answer only BALL if the red ball is closer, or ARROW if the green arrow is closer, nothing else.’’ and the prompt for the VPT-basic task was ‘‘In this image, if viewed from the perspective of the green 3D arrow in the direction the arrow is pointing, can a human see the red ball? Answer only YES or NO, nothing else’’. We evaluated each model’s generated responses across multiple temperatures, ranging from $0.0$ to $0.7$ in increments of $0.1$, and we report the average of the best 3 runs. Note that while this evaluation approach gives the VLMs more opportunities to perform well on our benchmark than other models, they still struggled immensely (see main text).
A.7 VLM Chain-of-Thought Evaluation
Chain-of-Thought (COT) prompting is a technique used to improve VLM and LLM performance by showing models the conceptual steps that might be helpful for solving a problem. COT has proven very effective on reasoning tasks. To understand if COT would also help DNNs solve VPT, we devised a prompting approach described below (see Fig. 7 for an example image used when prompting for COT reasoning). However, COT did not improve VLM accuracy on VPT (it achieved chance accuracy).
Figure 7: An example image for Chain-of-Thought prompting of vision language models to solve VPT.
Depth order: Here is a sample training image, from the perspective shown, is the red ball closer to the observer or is the green arrow closer to the observer? Answer only ’BALL’ if the red ball is closer, or ’ARROW’ if the green arrow is closer, nothing else. To answer this, let’s think step by step: The red ball is above the suitcase, which is a small distance away from the observer, approximately 3 feet. The green arrow is in front of the suitcase, which is closer to the observer, approximately 1 foot away. Thus the answer is: ARROW.
VPT: Here is a sample training image, if viewed from the perspective of the green 3D arrow, in the direction the arrow is pointing, can a human see the red ball? Answer only ’YES’ or ’NO’, nothing else. Let’s think step by step: If a human were looking in the direction the arrow was pointing, they would face a few possible occlusions, such as the suitcase. The red ball is above the suitcase, so the suitcase wouldn’t block the view of the red ball. Thus, the answer is: YES
A.8 Stable Diffusion Evaluation
We followed the method of Li et al. (2023) to evaluate Stable Diffusion 2.0 on the 3D-PC. This involved trying multiple prompts to optimize the zero-shot classification performance of the Stable Diffusion 2.0 model, on VPT-basic and depth order tasks. For VPT-basic we found that the prompt "A photo with red ball is visible from the green arrow’s perspective" for positive class and "A photo with red ball not visible from the green arrow’s perspective" for the negative class led to the best performance. For the depth order task, the prompt with the highest performance was "A photo with green arrow closer to the camera as compared to red ball" and "A photo with red ball closer to the camera as compared to green arrow" for positive and negative classes respectively.
A.9 Human Benchmark
@@
A.9.1 Experiment design
At the outset of the experiment, we acquired participant consent through a form approved by the Institutional Review Board (IRB). The experiment was performed on a computer using the Chrome browser. Following consent, we presented a demonstration with instructions and an example video. Participants had the option to revisit the instructions at any time during the experiment by clicking a link in the top right corner of the navigation bar.
Figure 8: An experiment trial.
@@ A.9.1 Experiment design
A.10 Human vs. DNN decision making on VPT-basic
@@
We compared the decision strategies of humans and DNNs on VPT-basic by measuring the correlations between their error patterns with Cohen’s $\kappa$ (Geirhos et al., 2020b). Model $\kappa$ scores were mostly correlated with accuracy on VPT-basic after linear probes and fine-tuning (Fig. 9). However, while nearly all DNNs were highly correlated with human error patterns after fine-tuning, the correlation between $\kappa$ scores and task accuracy disappeared (Fig. 9B, purple dots).
@@
Figure 9: Error pattern correlations (Cohen’s $\kappa$) between humans and DNNs on VPT-basic and Depth order.
Figure 10: DNNs fine-tuned on VPT-basic learn to solve the task by overfitting to the positions of the target objects. (A, B, C) There are significant correlations between DNN accuracy on ImageNet, fine-tuned accuracy on Depth/VPT-basic tasks, and the extent to which their attribution maps localize the target objects in VPT. Given the dramatic failure of DNNs to solve VPT-Strategy, this finding implies that DNNs rely on a and brittle strategy of memorizing object features like size and position to solve VPT.
Figure 11: Examples of decision attribution maps from a ViT large. (A) Attribution maps from linear probed and fine-tuned DNNs on VPT-basic. (B) Attribution maps from linear probed and fine-tuned DNNs on depth order.